OpenStudy (anonymous):

solve for x 7x + 7=12x - 2

OpenStudy (anonymous):

7x + 7=12x - 2 7x-12x=-2-7 -5x=-9 x=-9/-5

OpenStudy (anonymous):

@cassytaylor92 you are really good at algebra thank you for all of the help your a genius lol

OpenStudy (anonymous):

my pleasure :)